Jessicakuck:

total amount = P (1 + i)t Wallace took out a $5,000 loan for six years. He is being charged 4 percent interest, compounded annually. Calculate the total amount he will pay. $6,326.60 $6,312.38 $6,077.53

SkyVoltage43:

A

SkyVoltage43:

i gtg bye

SkyVoltage43:

jay u mind checking my work

SkyVoltage43:

A=P(1+r/n)^nt A=5000(1+.04/1)^1*6 A=5000(1.04)^6 A=5000(1.265319018) A=6,326.60